# AVOCADO: Adaptive Optimal Collision Avoidance driven by Opinion

### v1

@@ IV-D Exploiting attention to overcome deadlocks
The previous sections address the problem of unknown degree of cooperation in collision avoidance. However, there is an additional issue, typical in geometrical methods like ORCA, that is characterized by navigation deadlocks under symmetrical configurations [75]. To overcome them, AVOCADO exploits attention $A_{i}$.
@@
Our solution consists in injecting some small noise to the perceived velocity of agent $i$ when the agent is far from the robot, decreasing the noise to zero when the robot and the agent are close to a potential collision. When the robot and the agent are far from each other, the impact of noise is negligible because attention $A_{i}\approx 0$ and bifurcation does not happen. This preserves the neutrality of the opinion on the (shifted) degree of cooperation. When the robot and the agent are close to each other, noise is removed, so the performance guarantees on the adaptive law and the collision avoidance are preserved.
@@
AVOCADO is summarized in Algorithm 1. By combining a geometrical approach with a nonlinear opinion dynamics adaptive law, AVOCADO is able to adapt in real-time to the unknown degree of cooperation of the agents in a multi-agent setting. No communication is involved, only relying on the perceived position and velocity of the nearby agents. All the mathematical operations are inexpensive to compute, except for the optimization program in line 11. Nonetheless, since it entails a linear program, with an efficient implementation it is proven that a robot can process thousands of nearby agents in milliseconds [11]. We validate the computational simplicity of AVOCADO in Section V.
@@ V-A Head-on scenarios
Figure 5: Head-on scenarios for different ratio $b_{i}/d_{i}$ and degree of cooperation. The robot (blue) navigates using AVOCADO, whereas the agent is either non-cooperative (black, left column) or navigates using AVOCADO (red, right column). The motion of the robot and the agent is represented with disks of increasing transparency as time advances, and solid dots represent the center of the disks.
We simulate head-on scenarios to analyze the impact of the different parameters in AVOCADO. A robot and either a cooperative robot or a non-cooperative agent face each other, where their goals are the starting position of the other. Fig. 5 illustrates these scenarios for different values of bias $b_{i}$ (a, c, e for the non-cooperative agent; b, d, f for the cooperative robot). The robot using AVOCADO starts, in scenarios a) and b), assuming that the other robot or agent is non-cooperative with a low degree of cooperation; in scenarios c) and d), the robot initially assumes reciprocity, with a degree of cooperation equal to $0.5$; and in e) and f) the robot initially assumes a great degree of cooperation. Extreme bias values can lead to collisions, as in b) and e). In a), the robot smoothly avoids the collision with the agent, as the highly negative biased initial degree of cooperation (pessimistic robot) corresponds to reality. However, when the bias is incorrect, even if the agent is cooperative as in b), the robot takes some time to recover from the incorrect prior, resulting in shaky trajectories and even collisions in b). The opposite conflict arises when there is an initial highly positive bias in the degree of cooperation (optimistic). In case e), the robot is unable to correct the estimated degree of cooperation fast enough to realize that the agent is non-cooperative and take all the responsibility to avoid collision. Collision is avoided in case f), but the extreme bias leads to sub-optimal trajectories. Finally, a balanced bias allows the robot to react effectively to non-cooperative agents in c) and cooperative agents in d).
The impact of the other parameters is evaluated by studying the evolution of $x_{i}$ when the parameters change. The default parameters are specified in Table I. Now, we focus on the non-cooperative head-on scenario.
@@
Fig. 6 depicts the evolution of $x_{i}$ for the different cases. When the simulation starts, $x_{i}=b_{i}/d_{i}$. Then, as the robot detects that agent $i$ does not cooperate, $x_{i}$ decreases, trying to reach $x_{i}=-1$. Finally, when the collision is avoided and $A_{i}=0$, $x_{i}=b_{i}/d_{i}$ again. In some cases, as in Fig. 6e) with $\delta=0.1$, $x_{i}$ does not recover to $x_{i}=b_{i}/d_{i}$ because agent $i$ leaves the sensor range of the robot before letting the nonlinear opinion dynamics to converge to $x_{i}=b_{i}/d_{i}$. First, regarding $d_{i}$ (Eq. (10)), greater values imply a greater forgetting factor and therefore, a faster convergence according to the eigenvalue study from Eq. (16). For very large values ($d_{i}=6.5$), the nonlinear opinion dynamics are too reactive to the estimate $y_{i}$ and to the noise introduced (Eq. 19), leading to abrupt changes in $x_{i}$. Second, regarding $a_{i},c_{i}$ (Eq. (10)), since $c_{i}=1-a_{i}$ and $a_{i}\in\{0.1,0.3,0.6,0.9\}$, greater values of $a_{i}$ imply less reactivity against changes in the estimate $y_{i}$ and, henceforth, slowest convergence of the nonlinear opinion dynamics. Third, regarding $\varepsilon$ (Eq. (15)), it is observed that beyond some value around $\varepsilon=1$, the estimate of $y_{i}$ is reactive enough to guarantee a fast convergence of the nonlinear opinion dynamics. Fourth, the behavior of $\kappa_{i}$ (Eq. (11)) is similar to that of $\varepsilon$, in the sense that there is a point beyond which increasing $\kappa$ does not provide any improvement. Fifth, the same holds for $\delta_{i}$ (Eq. (11)). Finally, regarding the ratio $b_{i}/d_{i}$ (Eq. (10)), it is very interesting to observe that different values lead to different initial and final equilibrium, equal to $x_{i}=b_{i}/d_{i}$. In all the cases, since the agent is always non-cooperative, $x_{i}$ evolves towards a low value in order to take the responsibility of avoiding the collision. The plots of Fig. 6 show that, in general, very low parameters’ values are not desirable, as they make the opinion grow slowly and the robot to react late. Nevertheless, very large values can make the nonlinear opinion dynamics adaptive law to be too sensitive to changes in the agent and the estimate $y_{i}$, leading to a non-smooth evolution of $x_{i}$. Therefore, it is suggested to design the parameters as low as possible to guarantee collision avoidance but obtain smooth maneuvers. The nominal parameters described in Table I have been tuned using Bayesian Optimization, constrained the domain of the parameters according to the conclusions drawn from Fig. 6.
@@ V-B Multi-agent scenarios
We show the qualitative behavior of the robots in non-cooperative circle and crossing scenarios in Fig. 7. The planners address the navigation conflict at the center of the circle in two different ways. T-MPC and RVO-RL completely avoid dangerous zones by taking a detour; this might not be possible in a constrained scenario with boundaries. Among the other three approaches, ORCA is the one that takes more risks, since it assumes reciprocity in the degree of cooperation, colliding in both scenarios. SARL and AVOCADO achieve adaptation to the unknown degree of cooperation of the agents and other robots and manage to reach the goal in both scenarios; nonetheless, the movements exerted by the robots using AVOCADO are much smoother than those exerted by SARL, which is a key aspect when transferring navigation policies to real robots.
@@
Fig. 8 shows examples of scenarios only populated with robots. Fig. 8a demonstrates the deadlocks suffered by reciprocal approaches like ORCA when symmetries are encountered. AVOCADO overcomes this problem, as seen in Fig. 8e, by exploiting the attention mechanism in the nonlinear opinion dynamics adaptive law. T-MPC and RVO-RL take long unnecessary detours, as they are designed for scenarios with fewer robots and to take larger safety margins. T-MPC even presents many collisions, as it is not prepared to interact with other robots, and some of the RVO-RL robots do not reach the goal before the time out of $100$s. Trajectories of robots using SARL lack, again, smoothness, probably due to a reciprocal dance problem, specially noticeable in circle scenarios where this leads to collisions. All the robots using AVOCADO reach their goals avoiding collisions, with short path lengths and smooth trajectories, in both scenarios.
@@
Figure 8: Cooperative circle (top) and crossing (bottom) scenarios with 10 robots and different planners. Each robot is depicted with a different color. The representation follows the same rules in Fig. 7.
@@
Figure 9: Mixed cooperative/non-cooperative circle (top) and crossing (bottom) scenario with 4 agents, 6 robots and different planners. Each robot is depicted with a different color. The representation follows the same rules in Fig. 7.
@@
We define the success rate as a evaluation metric. Let $\mathsf{M}$ be the number of runs; $\mathrm{succ}_{j,k}\in\{0,1\}$ the indicator that is equal to $1$ if $||\mathbf{p}_{r,j}-\mathbf{p}_{r,j}^{*}||<\xi$ at the end of run $k$ or 00 otherwise, where $\mathbf{p}_{r,j}$ is the position of robot $j$, $\mathbf{p}^{*}_{r,j}$ is the desired goal of robot $j$, and $\xi>0$ is a small tolerance. Then, the success rate is defined as
@@
The success rates are in Fig. 10 and Fig. 11 for circle and crossing scenarios, respectively. AVOCADO outperforms all other approaches in success rate. AVOCADO_2, with a higher value of $d_{i}$ than AVOCADO_2, has a greater success rate in circle scenarios where the number of cooperative robots is greater, as it is more sensitive to the introduced noise that breaks symmetries. AVOCADO_1 is more stable and has a better performance in partially cooperative scenarios. AVOCADO_3 presents a worse success rate due to the bias, that assumes low degrees of cooperation when it is often not true. In this sense, it is very interesting to see how AVOCADO_4, with a bias towards great degrees of cooperation, presents the worst results among the AVOCADO versions in Fig. 10 unless in fully-cooperative cases. However, AVOCADO_4 achieves the best success rates in Fig.11. This is probably due to the fact that, in crossing scenarios, the robot first faces non-cooperative perpendicular traffic, where being cautious is desirable; meanwhile, in circle scenarios the robot faces agents coming in their same direction, leading to conflicts similar to the one previously seen in Fig. 5b. ORCA experiences deadlocks from geometrical symmetries, so it fails in all cooperative circle scenarios. The performance of T-MPC degrades when the number of cooperative agents increases, since, as it is observed in [65, 22], the method is suited for scenarios with less than 10 robots or agents. RVO-RL, tuned for fully-cooperative environments, achieves its best performance for $\mathsf{P}=1$, decreasing the success rate as non-cooperative agents appear. SARL, due to the non-smooth behavior, obtains poor success rates.
Figure 10: Success rates of robots with the different planners in circle scenarios. Red crosses mark the best planner.
Figure 11: Success rates of robots with the different planners in crossing scenarios.
@@
where $t_{j,k}>0$ denotes the instant when robot $j$ reaches its goal at run $k$. The results are aligned with the qualitative results depicted in Figs. 7, 8, and 9). RL-RVO and SARL exhibit the largest time to goal because their trajectories are irregular and present many detours. Meanwhile, the rest of the planners reach the goal in similar times. T-MPC manifests slightly lower times to reach the goal even tough the resulting trajectories are longer than those of AVOCADO or ORCA. This is due to the fact that T-MPC exerts higher velocities and it is, therefore, more risky, which explains its success rate metrics shown in Figs. 10 and 11.
@@ VI Experimental results
In Fig. 15a, pedestrians participating in the circle experiment walked with different velocities. When the person starting in the down-right leaves the intersection, the person starting in the top-center is still starting the trajectory. The robots adapt to the different velocities and safely reach the goal with smooth trajectories. The purple robot traverses the intersection moving to its right side, as there is free space there. Fig. 15b depicts a situation where the human starting in the down-right and the one starting in the top-center have similar behaviors as in the previous experiment in terms of speed. The pedestrian starting in the bottom-left, however, is faster than its previous homologous. The robots accomodate to this situation by being more cooperative and more reactive (yellow and blue) or choosing a different trajectory in the space that pedestrian is leaving behind (purple).
@@
Overall, extensive real experiments with multiple robots and humans prove that AVOCADO is effective navigation strategy for collision avoidance in mixed cooperative/non-cooperative environments. AVOCADO is a zero-shot approach since no further tuning is required to transfer the algorithm from simulations to real robots. In this sense, AVOCADO preserves the properties observed in simulations in terms of success rate, smoothness, adaptation to unknown degrees of cooperation and computational efficiency.

### v4

@@ IV-D Exploiting attention to avoid symmetry deadlocks
The previous sections address the problem of unknown degree of cooperation in collision avoidance. However, there is an additional issue, typical in VO-based methods like ORCA, that is characterized by navigation deadlocks under symmetrical configurations [86]. Among other mechanisms, VO-based methods usually solve symmetry deadlocks by injecting some arbitrary noise in the perceived velocity, where this amount is typically handcrafted. Instead, we propose an adaptive noise injection mechanism that exploits the attention $A_{i}$ of the opinion dynamics to correlate the amount of noise with safety.
@@
Specifically, our solution consists in injecting some small noise to the perceived velocity of agent $i$ when the agent is far from the robot, decreasing the noise to zero when the robot and the agent are close to a potential collision. When the robot and the agent are far from each other, the impact of noise is negligible because attention $A_{i}\approx 0$ and bifurcation does not happen. This preserves the neutrality of the opinion on the (shifted) degree of cooperation. When the robot and the agent are close to collision, noise is removed, so the performance guarantees on the adaptive law and the collision avoidance are preserved.
@@
AVOCADO is summarized in Algorithm 1. By combining a geometrical approach with a nonlinear opinion dynamics adaptive law, AVOCADO is able to adapt in real-time to the unknown degree of cooperation of the agents in a multi-agent setting. The unknown degree of cooperation of the agent induces an additional layer of complexity in collision avoidance, since, as it is shown in Section V, an incorrect assumption on the degree of cooperation leads to collision to existing geometrical, learning and model predictive approaches. On the other hand, AVOCADO is agnostic to the nature of the agents the robot encounters, so the robot can take a fast and flexible decision irrespective of the dynamic model of the agent. AVOCADO explicitly couples prediction with planning by estimating the effect its motion is producing in the other agents it encounters through the degree of cooperation $\alpha_{i}$ $\forall i\in\mathcal{N}$. In other words, the degree of cooperation $\alpha_{i}$, unless 00, expresses the fact that the motion of agent $i$ is influenced by the velocity exerted by the robot. By designing an adaptive law that estimates $\alpha_{i}$ in the continuous domain $\alpha_{i}\in[0,1]$, our method is able to deconflict the undesirable effects of such coupling by predicting how the agent reacts to the robot motion and plan the collision avoidance maneuver accordingly. No communication is involved, only relying on the perceived position and velocity of the nearby agents. All the mathematical operations are inexpensive to compute, except for the optimization program in line 11. Nonetheless, since it entails a linear program, with an efficient implementation it is proven that a robot can process thousands of nearby agents in milliseconds [12]. We validate the computational simplicity of AVOCADO in Section V.
@@ V-A Head-on scenarios
Figure 6: Head-on scenarios for different ratio $b_{i}/d_{i}$ and degree of cooperation. The robot (blue) navigates using AVOCADO, whereas the agent is either non-cooperative (black, left column) or navigates using AVOCADO (red, right column). The motion of the robot and the agent is represented with disks of increasing transparency as time advances, and solid dots represent the center of the disks. Collision is marked with crosses surrounded by circles
We simulate head-on scenarios to analyze the impact of the different parameters in AVOCADO. A robot and either a cooperative robot or a non-cooperative agent face each other, where their goals are the starting position of the other. Fig. 6 illustrates these scenarios for different values of bias $b_{i}$ (a, c, e for the non-cooperative agent; b, d, f for the cooperative robot). The robot using AVOCADO starts, in scenarios a) and b), assuming that the other robot or agent is non-cooperative with a low degree of cooperation; in scenarios c) and d), the robot initially assumes reciprocity, with a degree of cooperation equal to $0.5$; and in e) and f) the robot initially assumes a great degree of cooperation. Extreme bias values can lead to collisions, as in b) and e). In a), the robot effectively avoids the collision with the agent, as the highly negative biased initial degree of cooperation (pessimistic robot) corresponds to reality. However, when the bias is incorrect, even if the agent is cooperative as in b), the robot takes some time to recover from the incorrect prior, resulting in collision. The opposite conflict arises when there is an initial highly positive bias in the degree of cooperation (optimistic). In case e), the robot is unable to correct the estimated degree of cooperation fast enough to realize that the agent is non-cooperative and take all the responsibility to avoid collision. Collision is avoided in case f), but the extreme bias leads to sub-optimal trajectories. Finally, a balanced bias allows the robot to react effectively to non-cooperative agents in c) and cooperative agents in d).
The impact of the other parameters is evaluated by studying the evolution of ${o}_{i}$ when the parameters change. The default parameters are specified in Table I. Now, we focus on the non-cooperative head-on scenario.
@@
Fig. 7 depicts the evolution of ${o}_{i}$ for the different cases. As the robot detects that agent $i$ does not cooperate, ${o}_{i}$ decreases, trying to reach ${o}_{i}=-1$. Finally, when the collision is avoided and $A_{i}=0$, ${o}_{i}$ recovers to the initial value. In some cases, as in Fig. 7e) with $\delta=0.1$, ${o}_{i}$ does not recover to the initial value because agent $i$ leaves the sensor range of the robot before letting the nonlinear opinion dynamics to converge to ${o}_{i}=b_{i}/d_{i}$. First, regarding $d_{i}$ (Eq. (10)), greater values imply a greater forgetting factor and therefore, a faster convergence. For very large values ($d_{i}=6.5$), the nonlinear opinion dynamics are too reactive to the estimate ${e}_{i}$ and to the noise introduced (Eq. 19), leading to abrupt changes in ${o}_{i}$. Second, regarding $a_{i},c_{i}$ (Eq. (10)), since $c_{i}=1-a_{i}$ and $a_{i}\in\{0.1,0.3,0.6,0.9\}$, greater values of $a_{i}$ imply less reactivity against changes in the estimate ${e}_{i}$ and, henceforth, slowest convergence of the nonlinear opinion dynamics. Third, regarding $\varepsilon$ (Eq. (15)), it is observed that beyond some value around $\varepsilon=1$, the estimate of ${e}_{i}$ is reactive enough to guarantee a fast convergence of the nonlinear opinion dynamics. Fourth, the behavior of $\kappa_{i}$ (Eq. (11)) is similar to that of $\varepsilon$, in the sense that there is a point beyond which increasing $\kappa$ does not provide any improvement. Fifth, the same holds for $\delta_{i}$ (Eq. (11)). Finally, regarding the ratio $b_{i}/d_{i}$ (Eq. (10)), it is very interesting to observe that different values lead to different initial and final equilibrium, equal to ${o}_{i}=b_{i}/d_{i}$. In all the cases, since the agent is always non-cooperative, ${o}_{i}$ evolves towards a low value in order to take the responsibility of avoiding the collision. The plots of Fig. 7 show that, in general, very low parameters’ values are not desirable, as they make the opinion grow slowly and the robot to react late. Nevertheless, very large values can make the nonlinear opinion dynamics adaptive law to be too sensitive to changes in the agent and the estimate ${e}_{i}$, leading to a non-smooth evolution of ${o}_{i}$. Therefore, it is suggested to design the parameters as low as possible to guarantee collision avoidance but obtain fluid maneuvers. The nominal parameters described in Table I have been tuned using Bayesian Optimization, constrained the domain of the parameters according to the conclusions drawn from Fig. 7.
@@ V-B Multi-agent scenarios
We show the qualitative behavior of the robots in non-cooperative circle and crossing scenarios in Fig. 8. The planners address the navigation conflict at the center of the circle in two different ways. T-MPC and RVO-RL completely avoid dangerous zones by taking a detour; this might not be possible in a constrained scenario with boundaries. Among the other three approaches, ORCA is the one that takes more risks, since it assumes reciprocity in the degree of cooperation, colliding in both scenarios. SARL and AVOCADO achieve adaptation to the unknown degree of cooperation of the agents and other robots and manage to reach the goal in both scenarios; nonetheless, the movements exerted by the robots using AVOCADO are much more seamless than those exerted by SARL, which is a key aspect when transferring navigation policies to real robots.
@@
Fig. 9 shows examples of scenarios only populated with robots. Fig. 9a demonstrates the deadlocks suffered by reciprocal approaches like ORCA when symmetries are encountered. AVOCADO overcomes this problem, as seen in Fig. 9e, by exploiting the attention mechanism in the nonlinear opinion dynamics adaptive law. T-MPC and RVO-RL take long unnecessary detours, as they are designed for scenarios with fewer robots and to take larger safety margins. T-MPC even presents many collisions, as it is not prepared to interact with other robots, and some of the RVO-RL robots do not reach the goal before the time out of $100$s. Trajectories of robots using SARL are, again, very irregular, probably due to a reciprocal dance problem, specially noticeable in circle scenarios where this leads to collisions. All the robots using AVOCADO reach their goals avoiding collisions, with short path lengths and fluid trajectories, in both scenarios.
@@
Figure 9: Cooperative circle (top) and crossing (bottom) scenarios with 10 robots and different planners. Each robot is depicted with a different color. The representation follows the same rules in Fig. 8. Collisions are marked with crosses surrounded by circles
@@
Figure 10: Mixed cooperative/non-cooperative circle (top) and crossing (bottom) scenario with 4 agents, 6 robots and different planners. Each robot is depicted with a different color. The representation follows the same rules in Fig. 8. Collisions are marked with crosses surrounded by circles
@@
We define the success rate as a evaluation metric. Let $\mathsf{M}$ be the number of runs; $\mathrm{succ}_{j,k}\in\{0,1\}$ is the indicator that is equal to $1$ if $||\mathbf{p}_{r,j}-\mathbf{p}_{r,j}^{*}||<\xi$ at the end of run $k$ or 00 otherwise (it is also zero when the robot $j$ at run $k$ collides), where $\mathbf{p}_{r,j}$ is the position of robot $j$, $\mathbf{p}^{*}_{r,j}$ is the desired goal of robot $j$, and $\xi>0$ is a small tolerance. Then, the success rate is defined as
@@
The success rates are in Fig. 11 and Fig. 12 for circle and crossing scenarios, respectively. AVOCADO outperforms all other approaches in success rate. AVOCADO_2, with a higher value of $d_{i}$ than AVOCADO_1, has a greater success rate in circle scenarios where the number of cooperative robots is greater, as it is more sensitive to the introduced noise that breaks symmetries. AVOCADO_1 is more stable and has a better performance in partially cooperative scenarios. AVOCADO_3 presents a worse success rate due to the bias, that assumes low degrees of cooperation when it is often not true. In this sense, it is very interesting to see how AVOCADO_4, with a bias towards great degrees of cooperation, presents the worst results among the AVOCADO versions in Fig. 11 unless in fully-cooperative cases. However, AVOCADO_4 achieves the best success rates in Fig.12. This is probably due to the fact that, in crossing scenarios, the robot first faces non-cooperative perpendicular traffic, where being cautious is desirable; meanwhile, in circle scenarios the robot faces agents coming in their same direction, leading to conflicts similar to the one previously seen in Fig. 6b. ORCA experiences deadlocks from geometrical symmetries, so it fails in all cooperative circle scenarios. The performance of T-MPC degrades when the number of cooperative agents increases, since, as it is observed in [68, 23], the method is suited for scenarios with less than 10 robots or agents. RVO-RL, tuned for fully-cooperative environments, achieves its best performance for $\mathsf{P}=1$, decreasing the success rate as non-cooperative agents appear. SARL, due to the non-fluid behavior, obtains poor success rates.
Figure 11: Success rates of robots with the different planners in circle scenarios. Red crosses mark the best planner. The color bar refers to success rate as defined in Section V-B, whereas $\mathsf{P}$ refers to the proportion of cooperative robots. The yellower the better.
Figure 12: Success rates of robots with the different planners in crossing scenarios. The color bar refers to success rate as defined in Section V-B, whereas $\mathsf{P}$ refers to the proportion of cooperative robots. The yellower the better.
@@
where $t_{j,k}>0$ denotes the instant when robot $j$ reaches its goal at run $k$. The results are aligned with the qualitative results depicted in Figs. 8, 9, and 10). RL-RVO and SARL exhibit the largest time to goal because their trajectories are irregular and present many detours. Meanwhile, the rest of the planners reach the goal in similar times. T-MPC manifests slightly lower times to reach the goal even tough the resulting trajectories are longer than those of AVOCADO or ORCA. This is due to the fact that T-MPC exerts higher velocities and it is, therefore, more risky, which explains its success rate metrics shown in Figs. 11 and 12.
@@
V-C Static obstacles
As real-world scenarios often include other kind of obstacles, we study the behavior of AVOCADO in the presence of static obstacles. We followed the same approach of previous VO-based methods [28, 12, 35, 29, 13], which consider representing the boundaries of the obstacles as segments. Half-spaces constraining the velocities that lead to collision with these segments are directly included in the linear problem optimization. In this way, collision avoidance with static obstacles is naturally achieved with no extra cost.
To verify this insight, we conduct simulated experiments where robots following AVOCADO navigate in a circular scenario that includes static obstacles. Fig. 15 shows that the robots are able to adjust their trajectories to obstacles of different shapes.
a) b)
Refer to caption Refer to caption
Figure 15: Multi-robot experiments with static obstacles.
Nevertheless, as previously stated, an important remark is that AVOCADO is a planner that locally selects a non-colliding velocity. Therefore, an additional high-level planner would be needed to navigate in the presence of non-convex obstacles, such as in a maze-like environment.
@@ VI Experimental results
In Fig. 17a, pedestrians participating in the circle experiment walked with different velocities. When the person starting in the down-right leaves the intersection, the person starting in the top-center is still starting the trajectory. The robots adapt to the different velocities and safely reach the goal with fluid trajectories. The purple robot traverses the intersection moving to its right side, as there is free space there. Fig. 17b depicts a situation where the human starting in the down-right and the one starting in the top-center have similar behaviors as in the previous experiment in terms of speed. The pedestrian starting in the bottom-left, however, is faster than its previous homologous. The robots accomodate to this situation by being more cooperative and more reactive (yellow and blue) or choosing a different trajectory in the space that pedestrian is leaving behind (purple).
@@
Overall, extensive hardware experiments with multiple robots and humans prove that AVOCADO is effective navigation strategy for collision avoidance in mixed cooperative/non-cooperative environments. AVOCADO is a zero-shot approach since no further tuning is required to transfer the algorithm from simulations to real robots. In this sense, AVOCADO preserves the properties observed in simulations in terms of success rate, fluidity, adaptation to unknown degrees of cooperation and computational efficiency.